# Spectral Masking and Interpolation Attack (SMIA): A Black-box Adversarial Attack against Voice Authentication and Anti-Spoofing Systems

### v3

@@ 4.1. Datasets
We selected two standard datasets to facilitate a comprehensive evaluation of our system under different conditions: ASVSpoof 2019 (Todisco2019ASVspoof) [^1] https://www.asvspoof.org for assessing spoofing countermeasures and LibriSpeech (panayotov2015librispeech) [^2] https://www.openslr.org/12 for end-to-end system evaluation.
@@ 4.2. Authentication Models
  • X-Vectors (xvectors): We use the popular X-vector architecture, using the well-regarded SpeechBrain toolkit (ravanelli2021speechbrain) [^3] https://speechbrain.readthedocs.io/. X-vectors are embeddings extracted from a Time Delay Neural Network (TDNN), which aggregates frame-level features into a single, fixed-dimensional, utterance-level representation. Their widespread adoption in both research and industry makes them a critical SOTA target for evaluation.
  • DeepSpeaker (deepspeaker) [^4] https://github.com/philipperemy/deep-speaker: This is a well-established open-source model that utilizes deep residual networks (ResNet) to learn discriminative speaker embeddings. Its architecture, often trained with triplet loss functions, has been influential in the field and serves as a robust SOTA baseline for academic comparisons.
  • Microsoft Azure SV (microsoft_speaker) [^5] https://learn.microsoft.com/en-us/azure/ai-services/speech-service/speaker-recognition-overview: To demonstrate the practical relevance of our attack, we target the Microsoft Azure Speaker Verification service. As a commercial, black-box system, its internal architecture is not public. Successfully challenging this model highlights the effectiveness of our methods against systems deployed in real-world applications.
@@ 4.3. Anti-Spoofing Countermeasures
We evaluate our attack against three anti-spoofing CMs. These models were specifically chosen as they are the top-performing and winning systems from the ASVspoof 2019 Challenge (Todisco2019ASVspoof). A key factor in their success is their ability to operate directly on raw audio waveforms, allowing them to capture subtle spoofing artifacts that are often lost in traditional feature-extraction pipelines (e.g., spectrograms).
  • RawNet2 (rawnet2) [^6] https://github.com/eurecom-asp/rawnet2-antispoofing: This model is a highly influential SOTA architecture that demonstrated excellent performance on the ASVspoof 2019 dataset (Todisco2019ASVspoof). It employs a combination of 1D convolutional layers and Gated Recurrent Units (GRUs) to effectively learn discriminative features from raw audio, establishing it as a formidable benchmark for competition.
  • RawGAT-ST (rawgat) [^7] https://github.com/eurecom-asp/RawGAT-ST-antispoofing: An advanced architecture that builds upon the success of similar raw-waveform models in the challenge. It integrates a Graph Attention Network (GAT) with self-attention mechanisms, enabling it to weigh the importance of different audio segments and achieve a leading performance on the ASVspoof 2019 dataset (Todisco2019ASVspoof).
  • RawPC-DARTS (rawdarts) [^8] https://github.com/eurecom-asp/pc-darts-anti-spoofing: This model represents the cutting edge in CM design, utilizing Neural Architecture Search (NAS) to discover a topology optimized for the spoofing detection task. This automated design approach yielded a model that achieved a state-of-the-art result on the ASVspoof 2019 benchmark (Todisco2019ASVspoof), making it a critical system to evaluate against.
@@ 5. Evaluation Methodology
    End-to-End System Evaluation on ASVspoof: To create a robust evaluation subset, we randomly selected 30 speakers from the ASVspoof 2019 LA Evaluation partition (Todisco2019ASVspoof). For each selected speaker, we used 5 of their genuine bona fide audio samples for enrollment in the VAS. We then randomly selected 5 of their corresponding pre-existing spoofed audio files to serve as the base samples for our SMIA attack, resulting in 150 attack trials (30 speakers $\times$ 5 samples).
@@
    End-to-End System Evaluation on LibriSpeech: Next, we evaluate the attack on a complete, end-to-end pipeline using the LibriSpeech dataset (panayotov2015librispeech). For each speaker, we randomly select a total of ten audio files. Five of these files are used for the speaker’s enrollment in the voice authentication system. The other five files, each paired with a unique text prompt, are used to generate five initial spoofed samples using the Fish Speech APIs (fish-speech-v1.4). These baseline spoofed samples are then fed into our proposed SMIA attack to generate the final adversarial audios. It is these adversarial audios that are subsequently tested against the combined CM and VAS systems to simulate a full attack scenario.
@@
      Benchmark Attacks on Anti-Spoofing Systems (ASVspoof): To ensure a fair comparison with prior work (kassis2023breaking; hai2023sifdetectcracker), we applied our SMIA framework to spoofed audios from the ASVspoof 2019 dataset (Todisco2019ASVspoof). This methodology allows for a direct performance benchmark against existing methods. Specifically, we targeted all 13 logical access spoofing attack types (A07-A19) present in the evaluation partition. For each attack type, we randomly selected 50 audio samples, creating a comprehensive test set of 650 spoofed utterances (13 types $\times$ 50 samples) to ensure our attack’s generalizability.
@@
      Attacks on Anti-Spoofing Systems (LibriSpeech): We conducted a dedicated experiment targeting CMs. Test audios were generated using LibriSpeech (panayotov2015librispeech) and Fish Speech APIs (fish-speech-v1.4), then processed by our SMIA attack to test the vulnerability of various CMs.
@@
      Attacks on Voice Authentication Systems (VAS): We extensively evaluated VAS vulnerabilities using audio generated from LibriSpeech (panayotov2015librispeech) and Fish Speech (fish-speech-v1.4). The attacks were tested against VAS models across three biometric tasks: Speaker Verification (SV), Open-Set Identification (OSI), and Closed-Set Identification (CSI).
@@
      Computational Efficiency Evaluation: We evaluated SMIA’s efficiency against SiFDetectCracker (hai2023sifdetectcracker), measuring the iterations and time required for a successful attack.
@@
    Stealth and Detectability Analysis: To demonstrate that our attack is significantly harder to detect than prior work, we perform a comparative visual analysis of the audio spectrograms. We compare the adversarial samples generated by SMIA against those from SiFDetectCracker (hai2023sifdetectcracker) to show that our method avoids the structured, conspicuous artifacts that create a detectable pattern. This highlights our method’s superior stealth and resilience against basic forensic analysis.
@@ 6.1. End-to-End System Evaluation on ASVspoof
Our evaluation shows that the proposed method significantly outperforms Kassis et al. (kassis2023breaking). In contrast, a direct comparison with Hai et al. (hai2023sifdetectcracker) remains unfeasible, as their work does not address the combined VAS and CM systems utilized in our analysis. Consequently, this experiment provides evidence of superior performance where comparison is possible, while also providing a clear benchmark for future research on this specific dataset configuration.
@@ 6.2. End-to-End System Evaluation on LibriSpeech
For the more practical scenario using the LibriSpeech dataset (panayotov2015librispeech), Table 4 summarizes the end-to-end success rate across all nine combinations of CM and VAS models. The findings are exceptional, with the SMIA attack achieving a perfect 100% ASR in the majority of configurations. These results highlight the robustness of SMIA, demonstrating its ability to consistently bypass diverse end-to-end defenses.
@@ 6.3.1. Benchmark Attacks on Anti-Spoofing Systems using ASVspoof 2019
For a direct and fair comparison with prior work, our first targeted experiment benchmarked the SMIA framework on the ASVspoof 2019 dataset (Todisco2019ASVspoof). The results, presented in Table 5, are unequivocal. SMIA substantially outperforms both SOTA baselines across all tested countermeasures, demonstrating a significant advancement in attack efficacy.
@@
SiFDetectCracker (hai2023sifdetectcracker) 80.4 75.8 84.1
Kassis et al. (kassis2023breaking) 12.03 15.82
@@ 6.3.3. Attacks on Voice Authentication Systems
Our extensive evaluation of Voice Authentication Systems demonstrated their profound susceptibility to the SMIA attack. The results, detailed in Table 7, show the attack’s performance across Speaker Verification (SV), Open-Set Identification (OSI), and Closed-Set Identification (CSI) tasks. The key finding is the attack’s universal effectiveness; SMIA achieved an ASR of at least 87% across all tested systems, including DeepSpeaker (deepspeaker), X-Vectors (xvectors), and the commercial Microsoft Azure SV (microsoft_speaker), peaking at a perfect 100%.
@@
Kassis et al. (kassis2023breaking) SV X-Vectors - 100 - -
@@ 6.3.4. Computational Efficiency
For a holistic evaluation, we compared the efficiency of SMIA against the SiFDetectCracker baseline (hai2023sifdetectcracker) when targeting standalone CMs. We evaluated efficiency using two key metrics: (1) the average number of model queries required for a successful attack, and (2) the average computational time. The query count is a more robust, hardware-independent metric for BBO attacks, while the time cost demonstrates practical feasibility.
@@
Table 8. Efficiency comparison (Avg. Queries and Time) of SMIA and SiFDetectCracker (hai2023sifdetectcracker) against standalone CMs.
@@
Figure 3. Visual comparison of attack stealth. The spectrograms show an original audio versus adversarial samples generated by SiFDetectCracker (Hai et al.) (hai2023sifdetectcracker) and our three SMIA modes for two different source audios (top and bottom rows). Note the conspicuous “silent areas” introduced by the SiFDetectCracker (hai2023sifdetectcracker) (3(b), 3(e)), which create a detectable pattern. In contrast, the perturbations from SMIA (3(c), 3(f)) are more subtle and randomly distributed, making them significantly harder to detect.
@@ 6.5. Stealth and Detectability Analysis
To assess the stealth of our attack, we conducted a visual analysis comparing the spectrograms of our adversarial samples against those from Hai et al. (hai2023sifdetectcracker), as visualized in Figure 3. The spectrograms for the Hai et al. attack (Figures 3(b) and 3(e)) reveal a critical vulnerability: the method introduces highly structured and conspicuous artifacts in the form of clear “silent areas”. These predictable, zeroed-out regions are easily detectable by simple forensic analysis, and if filtered out, the attack’s effectiveness would drop dramatically. In contrast, all modes of our SMIA attack produce more subtle and randomly distributed perturbations that blend into the natural texture of the spectrogram. By avoiding obvious, structured patterns, SMIA generates adversarial samples that are inherently stealthier and more resilient to detection and removal by forensic countermeasures.
@@ 7.1. Ablation Study on Attack Modes
Finally, to understand the contribution of our different perturbation strategies, we conducted an ablation study comparing the Masking, Interpolation, and Hybrid modes. As shown in Table 10, the results highlight that the effectiveness of each mode is highly dependent on the target system’s composition. While the Interpolation mode appears dominant against RawNet2 (rawnet2) and RawGAT-ST(rawgat) (achieving 100% ASR with X-Vectors), it exhibits a critical vulnerability against RawPC-Darts (rawdarts), where its performance collapses to just 5%. Conversely, the Hybrid mode proves to be the most resilient overall. It sacrifices marginal performance on the simpler defenses to achieve a decisive 84.4% ASR against RawPC-Darts (rawdarts), demonstrating that combining strategies is essential for bypassing the most robust defense combinations.
@@ 7.2. Ablation Study on Randomness
To analyze the impact of our stochastic perturbation strategy, we conducted an ablation study comparing two of our proposed operational modes: a deterministic Simple mode and a stochastic Random mode. The results in Table 11 reveal a crucial insight: while the deterministic approach is highly potent against most systems, introducing randomness provides a decisive performance advantage against specific advanced defenses. Notably, against the DeepSpeaker (deepspeaker) and RawPC-DARTS combination, the Simple mode is effectively blocked (14% ASR). However, the Random mode recovers significant capability, boosting the ASR from 14% to 66.5%, highlighting the strategic value of stochasticity in overcoming robust countermeasures.
@@ 8. Discussion
SMIA represents a significant advancement in understanding the vulnerabilities of modern voice biometric security systems. The findings carry several critical implications for the field. The practical success of SMIA, achieving up to a 100% ASR against standalone countermeasures and at least 82.7% against combined, layered systems, underscores a fundamental flaw in current defense paradigms. The evaluation across a variety of models, including open-source standards like X-Vectors (xvectors) and commercial systems such as Microsoft Azure SV (microsoft_speaker), demonstrates the broad, real-world applicability of this attack methodology. A key strength of SMIA is its generalization; by successfully utilizing spoofed audio from the 13 different attack types in the ASVspoof 2019 dataset, this work has demonstrated that SMIA does not depend on the artifacts of a single voice synthesis method. This flexibility positions it as a persistent and evolving threat.

### v4

@@ 4.1. Datasets
We selected two standard datasets to facilitate a comprehensive evaluation of our system under different conditions: ASVSpoof 2019 (et al., 2019) [^1] https://www.asvspoof.org for assessing spoofing countermeasures and LibriSpeech (Panayotov et al., 2015) [^2] https://www.openslr.org/12 for end-to-end system evaluation.
@@ 4.2. Authentication Models
  • X-Vectors (Snyder et al., 2018): We use the popular X-vector architecture, using the well-regarded SpeechBrain toolkit (Ravanelli et al., 2021) [^3] https://speechbrain.readthedocs.io/. X-vectors are embeddings extracted from a Time Delay Neural Network (TDNN), which aggregates frame-level features into a single, fixed-dimensional, utterance-level representation. Their widespread adoption in both research and industry makes them a critical SOTA target for evaluation.
  • DeepSpeaker (Li et al., 2017) [^4] https://github.com/philipperemy/deep-speaker: This is a well-established open-source model that utilizes deep residual networks (ResNet) to learn discriminative speaker embeddings. Its architecture, often trained with triplet loss functions, has been influential in the field and serves as a robust SOTA baseline for academic comparisons.
  • Microsoft Azure SV (Microsoft, ) [^5] https://learn.microsoft.com/en-us/azure/ai-services/speech-service/speaker-recognition-overview: To demonstrate the practical relevance of our attack, we target the Microsoft Azure Speaker Verification service. As a commercial, black-box system, its internal architecture is not public. Successfully challenging this model highlights the effectiveness of our methods against systems deployed in real-world applications.
@@ 4.3. Anti-Spoofing Countermeasures
We evaluate our attack against three anti-spoofing CMs. These models were specifically chosen as they are the top-performing and winning systems from the ASVspoof 2019 Challenge (et al., 2019). A key factor in their success is their ability to operate directly on raw audio waveforms, allowing them to capture subtle spoofing artifacts that are often lost in traditional feature-extraction pipelines (e.g., spectrograms).
  • RawNet2 (Tak et al., 2021b) [^6] https://github.com/eurecom-asp/rawnet2-antispoofing: This model is a highly influential SOTA architecture that demonstrated excellent performance on the ASVspoof 2019 dataset (et al., 2019). It employs a combination of 1D convolutional layers and Gated Recurrent Units (GRUs) to effectively learn discriminative features from raw audio, establishing it as a formidable benchmark for competition.
  • RawGAT-ST (Tak et al., 2021a) [^7] https://github.com/eurecom-asp/RawGAT-ST-antispoofing: An advanced architecture that builds upon the success of similar raw-waveform models in the challenge. It integrates a Graph Attention Network (GAT) with self-attention mechanisms, enabling it to weigh the importance of different audio segments and achieve a leading performance on the ASVspoof 2019 dataset (et al., 2019).
  • RawPC-DARTS (Ge et al., 2021) [^8] https://github.com/eurecom-asp/pc-darts-anti-spoofing: This model represents the cutting edge in CM design, utilizing Neural Architecture Search (NAS) to discover a topology optimized for the spoofing detection task. This automated design approach yielded a model that achieved a state-of-the-art result on the ASVspoof 2019 benchmark (et al., 2019), making it a critical system to evaluate against.
@@ 5. Evaluation Methodology
    End-to-End System Evaluation on ASVspoof: To create a robust evaluation subset, we randomly selected 30 speakers from the ASVspoof 2019 LA Evaluation partition (et al., 2019). For each selected speaker, we used 5 of their genuine bona fide audio samples for enrollment in the VAS. We then randomly selected 5 of their corresponding pre-existing spoofed audio files to serve as the base samples for our SMIA attack, resulting in 150 attack trials (30 speakers $\times$ 5 samples).
@@
    End-to-End System Evaluation on LibriSpeech: Next, we evaluate the attack on a complete, end-to-end pipeline using the LibriSpeech dataset (Panayotov et al., 2015). For each speaker, we randomly select a total of ten audio files. Five of these files are used for the speaker’s enrollment in the voice authentication system. The other five files, each paired with a unique text prompt, are used to generate five initial spoofed samples using the Fish Speech APIs (Liao et al., 2024). These baseline spoofed samples are then fed into our proposed SMIA attack to generate the final adversarial audios. It is these adversarial audios that are subsequently tested against the combined CM and VAS systems to simulate a full attack scenario.
@@
      Benchmark Attacks on Anti-Spoofing Systems (ASVspoof): To ensure a fair comparison with prior work (Kassis and Hengartner, 2023; Hai et al., 2023), we applied our SMIA framework to spoofed audios from the ASVspoof 2019 dataset (et al., 2019). This methodology allows for a direct performance benchmark against existing methods. Specifically, we targeted all 13 logical access spoofing attack types (A07-A19) present in the evaluation partition. For each attack type, we randomly selected 50 audio samples, creating a comprehensive test set of 650 spoofed utterances (13 types $\times$ 50 samples) to ensure our attack’s generalizability.
@@
      Attacks on Anti-Spoofing Systems (LibriSpeech): We conducted a dedicated experiment targeting CMs. Test audios were generated using LibriSpeech (Panayotov et al., 2015) and Fish Speech APIs (Liao et al., 2024), then processed by our SMIA attack to test the vulnerability of various CMs.
@@
      Attacks on Voice Authentication Systems (VAS): We extensively evaluated VAS vulnerabilities using audio generated from LibriSpeech (Panayotov et al., 2015) and Fish Speech (Liao et al., 2024). The attacks were tested against VAS models across three biometric tasks: Speaker Verification (SV), Open-Set Identification (OSI), and Closed-Set Identification (CSI).
@@
      Computational Efficiency Evaluation: We evaluated SMIA’s efficiency against SiFDetectCracker (Hai et al., 2023), measuring the iterations and time required for a successful attack.
@@
    Stealth and Detectability Analysis: To demonstrate that our attack is significantly harder to detect than prior work, we perform a comparative visual analysis of the audio spectrograms. We compare the adversarial samples generated by SMIA against those from SiFDetectCracker (Hai et al., 2023) to show that our method avoids the structured, conspicuous artifacts that create a detectable pattern. This highlights our method’s superior stealth and resilience against basic forensic analysis.
@@ 6.1. End-to-End System Evaluation on ASVspoof
Our evaluation shows that the proposed method significantly outperforms Kassis et al. (Kassis and Hengartner, 2023). In contrast, a direct comparison with Hai et al. (Hai et al., 2023) remains unfeasible, as their work does not address the combined VAS and CM systems utilized in our analysis. Consequently, this experiment provides evidence of superior performance where comparison is possible, while also providing a clear benchmark for future research on this specific dataset configuration.
@@ 6.2. End-to-End System Evaluation on LibriSpeech
For the more practical scenario using the LibriSpeech dataset (Panayotov et al., 2015), Table 4 summarizes the end-to-end success rate across all nine combinations of CM and VAS models. The findings are exceptional, with the SMIA attack achieving a perfect 100% ASR in the majority of configurations. These results highlight the robustness of SMIA, demonstrating its ability to consistently bypass diverse end-to-end defenses.
@@ 6.3.1. Benchmark Attacks on Anti-Spoofing Systems using ASVspoof 2019
For a direct and fair comparison with prior work, our first targeted experiment benchmarked the SMIA framework on the ASVspoof 2019 dataset (et al., 2019). The results, presented in Table 5, are unequivocal. SMIA substantially outperforms both SOTA baselines across all tested countermeasures, demonstrating a significant advancement in attack efficacy.
@@
SiFDetectCracker (Hai et al., 2023) 80.4 75.8 84.1
Kassis et al. (Kassis and Hengartner, 2023) 12.03 15.82
@@ 6.3.3. Attacks on Voice Authentication Systems
Our extensive evaluation of Voice Authentication Systems demonstrated their profound susceptibility to the SMIA attack. The results, detailed in Table 7, show the attack’s performance across Speaker Verification (SV), Open-Set Identification (OSI), and Closed-Set Identification (CSI) tasks. The key finding is the attack’s universal effectiveness; SMIA achieved an ASR of at least 87% across all tested systems, including DeepSpeaker (Li et al., 2017), X-Vectors (Snyder et al., 2018), and the commercial Microsoft Azure SV (Microsoft, ), peaking at a perfect 100%.
@@
Kassis et al. (Kassis and Hengartner, 2023) SV X-Vectors - 100 - -
@@ 6.3.4. Computational Efficiency
For a holistic evaluation, we compared the efficiency of SMIA against the SiFDetectCracker baseline (Hai et al., 2023) when targeting standalone CMs. We evaluated efficiency using two key metrics: (1) the average number of model queries required for a successful attack, and (2) the average computational time. The query count is a more robust, hardware-independent metric for BBO attacks, while the time cost demonstrates practical feasibility.
@@
Table 8. Efficiency comparison (Avg. Queries and Time) of SMIA and SiFDetectCracker (Hai et al., 2023) against standalone CMs.
@@
Figure 3. Visual comparison of attack stealth. The spectrograms show an original audio versus adversarial samples generated by SiFDetectCracker (Hai et al.) (Hai et al., 2023) and our three SMIA modes for two different source audios (top and bottom rows). Note the conspicuous “silent areas” introduced by the SiFDetectCracker (Hai et al., 2023) (3(b), 3(e)), which create a detectable pattern. In contrast, the perturbations from SMIA (3(c), 3(f)) are more subtle and randomly distributed, making them significantly harder to detect.
@@ 6.5. Stealth and Detectability Analysis
To assess the stealth of our attack, we conducted a visual analysis comparing the spectrograms of our adversarial samples against those from Hai et al. (Hai et al., 2023), as visualized in Figure 3. The spectrograms for the Hai et al. attack (Figures 3(b) and 3(e)) reveal a critical vulnerability: the method introduces highly structured and conspicuous artifacts in the form of clear “silent areas”. These predictable, zeroed-out regions are easily detectable by simple forensic analysis, and if filtered out, the attack’s effectiveness would drop dramatically. In contrast, all modes of our SMIA attack produce more subtle and randomly distributed perturbations that blend into the natural texture of the spectrogram. By avoiding obvious, structured patterns, SMIA generates adversarial samples that are inherently stealthier and more resilient to detection and removal by forensic countermeasures.
@@ 7.1. Ablation Study on Attack Modes
Finally, to understand the contribution of our different perturbation strategies, we conducted an ablation study comparing the Masking, Interpolation, and Hybrid modes. As shown in Table 10, the results highlight that the effectiveness of each mode is highly dependent on the target system’s composition. While the Interpolation mode appears dominant against RawNet2 (Tak et al., 2021b) and RawGAT-ST(Tak et al., 2021a) (achieving 100% ASR with X-Vectors), it exhibits a critical vulnerability against RawPC-Darts (Ge et al., 2021), where its performance collapses to just 5%. Conversely, the Hybrid mode proves to be the most resilient overall. It sacrifices marginal performance on the simpler defenses to achieve a decisive 84.4% ASR against RawPC-Darts (Ge et al., 2021), demonstrating that combining strategies is essential for bypassing the most robust defense combinations.
@@ 7.2. Ablation Study on Randomness
To analyze the impact of our stochastic perturbation strategy, we conducted an ablation study comparing two of our proposed operational modes: a deterministic Simple mode and a stochastic Random mode. The results in Table 11 reveal a crucial insight: while the deterministic approach is highly potent against most systems, introducing randomness provides a decisive performance advantage against specific advanced defenses. Notably, against the DeepSpeaker (Li et al., 2017) and RawPC-DARTS combination, the Simple mode is effectively blocked (14% ASR). However, the Random mode recovers significant capability, boosting the ASR from 14% to 66.5%, highlighting the strategic value of stochasticity in overcoming robust countermeasures.
@@ 8. Discussion
SMIA represents a significant advancement in understanding the vulnerabilities of modern voice biometric security systems. The findings carry several critical implications for the field. The practical success of SMIA, achieving up to a 100% ASR against standalone countermeasures and at least 82.7% against combined, layered systems, underscores a fundamental flaw in current defense paradigms. The evaluation across a variety of models, including open-source standards like X-Vectors (Snyder et al., 2018) and commercial systems such as Microsoft Azure SV (Microsoft, ), demonstrates the broad, real-world applicability of this attack methodology. A key strength of SMIA is its generalization; by successfully utilizing spoofed audio from the 13 different attack types in the ASVspoof 2019 dataset, this work has demonstrated that SMIA does not depend on the artifacts of a single voice synthesis method. This flexibility positions it as a persistent and evolving threat.